OpenStudy (anonymous):

@hartnn can i please get some help you did it already with somebody else but could you help me please?

hartnn (hartnn):

whats your question ?

OpenStudy (anonymous):

Use the graph below to answer the question that follows:

OpenStudy (anonymous):

What is the rate of change between the interval x = π and x = three pi over two?

hartnn (hartnn):

know the rate of change formula ?

OpenStudy (anonymous):

But the thing is that my answer are... Pi/6 6/pi 4/pi pi/4

hartnn (hartnn):

yes, the correct answer is within those choices only, here's the formula \(\huge \dfrac{f(b)-f(a)}{b-a}\) here, a= \(\pi/2\) b= \(3\pi/2\)

OpenStudy (anonymous):

i saw how you did it before but the answer you got was 0.5 it's not the same as mine.. i know how to do the steps..

hartnn (hartnn):

lets do it :) numerator = f(3pi/2) - f(pi/2) know how to find f(3pi/2) and f(pi/2) ?

OpenStudy (anonymous):

i got the answer thank you....

OpenStudy (anonymous):

What about this one.. If sin Θ = negative square root 3 over 2 and π < Θ < 3 pi over 2, what are the values of cos Θ and tan Θ?

hartnn (hartnn):

if you want me to verify it, tell me what u got ? do you know the trigonometric relation, \(\Large \sin^2 x + \cos^2 x = 1\) ?

OpenStudy (anonymous):

yes....

hartnn (hartnn):

so you have value of sin theta, use that formula and try to get value of cos theta...try ?

OpenStudy (anonymous):

ok i will thank you..